### v2

# Helson’s conjecture for smooth numbers

Abstract.
Let $\Psi(x,y)$ denote the count of $y$-smooth numbers below $x$ and $P(n)$ denote the largest prime factor of $n$. We prove that for $f$ a Steinhaus random multiplicative function, the partial sums over $y$-smooth numbers enjoy better than squareroot cancellation, in the sense that
$\mathbb{E}\Big|\sum_{\begin{subarray}{c}1\leqslant n\leqslant x\\ P(n)\leqslant y\end{subarray}}f(n)\Big|=o\left(\Psi(x,y)^{1/2}\right),$
uniformly for $(\log x)^{30}\leqslant y\leqslant x$. Our bounds are quantitative and give a large saving when $y$ isn’t too close to $x$.
1. Introduction
1.1. Motivation and statement of results
Multiplicative functions are a fundamental object of study in number theory, and, in recent years, research on random multiplicative functions has been pursued as an avenue to better understand the behaviour of certain families of multiplicative functions. The study of random multiplicative functions was initiated by Wintner [Win] in 1944, where he introduced the Rademacher random multiplicative function as a model for studying the partial sums of the Möbius function. In this paper, we concern ourselves only with the Steinhaus random multiplicative function, $f\colon\mathbb{N}\rightarrow\mathbb{C}$, which is defined by letting $\bigl(f(p)\bigr)_{p\text{ prime}}$ be independent and identically distributed random variables uniformly distributed on the complex unit circle $\{z\in\mathbb{C}:|z|=1\}$, and extending to composite numbers $n=p_{1}^{\alpha_{1}}\dots p_{r}^{\alpha_{r}}$ by setting
$f(n)=\prod_{i=1}^{r}f(p_{i})^{\alpha_{i}}.$
The Steinhaus random multiplicative function serves as a model for families of Dirichlet characters, $n\rightarrow\chi(n)$ for $\chi\bmod r$ chosen uniformly at random, in addition to continuous characters, $n\rightarrow n^{it}$, for $t\in[T,2T]$ chosen uniformly at random, where $r$ and $T$ are to be thought of as parameters tending to infinity. For $f$ a Steinhaus random multiplicative function, we have “perfect orthogonality”, in the sense that
$\mathbb{E}\Bigl[f(n)\overline{f(m)}\Bigr]=\mathbf{1}_{n=m}.$
This relation is mirrored by corresponding orthogonality relations for characters. It is of great interest to analytic number theorists to understand the statistical behaviour of character sums, $\sum_{n\leqslant x}\chi(n)$, and zeta sums, $\sum_{n\leqslant x}n^{it}$. In light of the orthogonality relations, the second moments of these quantities are fairly straightforward to evaluate, most of all in the Steinhaus case where we immediately see that $\mathbb{E}\bigl|\sum_{n\leqslant x}f(n)\bigr|^{2}=\lfloor x\rfloor$. A natural problem that follows is to understand the first moment, $\mathbb{E}\bigl|\sum_{n\leqslant x}f(n)\bigr|$, in addition to the character and zeta sum analogues, though understanding these quantities is a much harder problem. A naive first guess is that, ignoring the multiplicative dependence structure, the sum $\sum_{n\leqslant x}f(n)$ may behave like the first absolute moment of a sum of mean-zero independent random variables with unit variance, and, if this were true, the Cauchy–Schwarz upper bound $\mathbb{E}\bigl|\sum_{n\leqslant x}f(n)\bigr|\leqslant\sqrt{x}$ would be sharp up to a constant. This upper bound is often referred to as squareroot cancellation, and more generally gives $\mathbb{E}\bigl|\sum_{n\in\mathcal{A}}f(n)\bigr|\leqslant\sqrt{|\mathcal{A}|}$ for an arbitrary set $\mathcal{A}$. Helson [Helson] conjectured that, for the full sum up to $x$, we should actually have more than squareroot cancellation, in the sense that $\mathbb{E}\bigl|\sum_{n\leqslant x}f(n)\bigr|=o(\sqrt{x})$. The answer to this conjecture remained unclear until the problem was fully resolved by Harper [HarperLow], who showed that Helson’s conjecture was true, proving the precise estimate
$\mathbb{E}\Bigl|\sum_{1\leqslant n\leqslant x}f(n)\Bigr|\asymp\frac{\sqrt{x}}{(\log\log x)^{1/4}}.$
The analogous character and zeta upper bounds are also proved by Harper in [harpertypical] using a remarkable derandomisation method. More precisely, Harper explicitly relates these moments of character sums and zeta sums to (a certain stage in the proof of) the random multiplicative case. As is to be expected, the length of the sum in that case depends on the parameters $r$ and $T$ that determine the size of the family of characters. For example, it is shown that for any large prime $r$ and for all $x\leqslant r$ with $x,\frac{r}{x}\to+\infty$, one has
$\frac{1}{r-1}\sum_{\chi\bmod r}\Big|\sum_{1\leqslant n\leqslant x}\chi(n)\Big|=o(\sqrt{x}).$
In that work, Harper also conjectured (see [harpertypical, Equation (1.2)]) that this phenomenon can be pushed further, in that the “conductor restriction” $x\leqslant r$ can be weakened if one has a Möbius twist, an application of which would be to break the classical “squareroot barrier” in Möbius cancellation (see [WXRatios] for further discussion). We shall discuss later in the introduction (after Theorem 1.1) about how an analogous conjecture in our setting may have consequences for counting smooth numbers in small intervals.
The key idea behind Harper’s theorem is to notice that the partial sums are connected to a phenomenon known as critical Gaussian multiplicative chaos (GMC). To briefly elaborate, we define $F_{x}(s)\coloneqq\prod_{p\leqslant x}\bigl(1-\frac{f(p)}{p^{s}}\bigr)^{-1}$ to be the truncated Euler product. By a non-trivial conditioning argument and an application of Plancherel’s identity, the proof works by first showing that
$\mathbb{E}\Bigl|\sum_{n\leqslant x}f(n)\Bigr|\approx\sqrt{\frac{x}{\log x}}\mathbb{E}\left(\int_{-1/2}^{1/2}|F_{x}(1/2+it)|^{2}dt\right)^{1/2},$ (1.1)
and then by showing that this expectation of the Euler product integral is smaller than naively anticipated. The driving force behind this is the fact that $\bigl(\log|F(1/2+it)|\bigr)_{t\in[-1/2,1/2]}$ is well approximated by a Gaussian field with logarithmic correlations, and the exponent 2 is exactly the critical value in this case, all together meaning that these integrals are precisely the setting where one observes critical Gaussian multiplicative chaos. The reader is encouraged to consult [Harperhigh, HarperIII] for more detailed discussions. We also remark that there have been recent progress in understanding the exact limiting distribution of the partial sums in [GW-Final, Hardy2025].
It is natural to ask whether one still obtains “better than squareroot cancellation” when the full sum is replaced by a sum over a set $\mathcal{A}$ that has interesting arithmetic structure. In previous work of the second author [Xu], the case where $\mathcal{A}=\mathcal{A}(x,R)$ is a set of $R$-rough numbers up to $x$, i.e., all elements in the set have prime factors with size at least $R$, was studied. In that case, there is an interesting transition range, since we observe “better than squareroot” cancellation,
$\mathbb{E}\Bigl|\sum_{n\in\mathcal{A}(x,R)}f(n)\Bigr|=o\left(\sqrt{|\mathcal{A}(x,R)|}\right),$
whenever $\log\log R$ is smaller than roughly $\sqrt{\log\log x}$, but the statement fails to hold when $R$ is any larger. The case where $\mathcal{A}$ is a short interval was studied in [Caichshort], and, together with the related work [Chang], these three papers observe some universality in the transition threshold which is ultimately related to the ballot problem.
In this paper, our attention lies on what happens when $\mathcal{A}$ is the set of $y$-smooth numbers, i.e., all elements in $\mathcal{A}$ only have prime factors $p\leqslant y$. Here and throughout the paper, let $P(n)$ denote the largest prime factor of $n$, so that $\Psi(x,y)=\#\{n\leqslant x:P(n)\leqslant y\}$ denotes the count of $y$-smooth numbers up to $x$. Note that the case where $y=x$ corresponds to the full sum, for which we know from Harper[HarperLow] that we have better than squareroot cancellation. Analogously to the previous cases, one might wonder if there is a certain transition range where the “better than squareroot” cancellation disappears. Our main theorem suggests that this may not be the case.
Theorem 1.1.
Suppose that $(\log x)^{30}\leqslant y\leqslant x$. Then we have
$\mathbb{E}\Big|\sum_{\begin{subarray}{c}1\leqslant n\leqslant x\\ P(n)\leqslant y\end{subarray}}f(n)\Big|=o\left(\Psi(x,y)^{1/2}\right),$
uniformly for $y$ in this range. Quantitative bounds can be found in Theorems 1.3 and LABEL:t:small_u_upper_bound.
One motivation and a potential application for our result is to make further progress on the classical problem of counting $y$-smooth numbers in short intervals $[x,x+h]$. See recent developments and discussions in [soundshortinterval, KYounis]. The breakthrough work of Matomäki and Radziwiłł [MR16] shows the existence of $x^{\varepsilon}$-smooth numbers unconditionally in intervals of length $h\gg_{\varepsilon}\sqrt{x}$, coming close to solving an old problem that, for any fixed $\varepsilon>0$, there exists an $x^{\varepsilon}$-smooth number in $x,x+\sqrt{x}]$ when $x$ is sufficiently large. To see how our result may be useful, similarly to [harpertypical, Equation (1.2)] (though removing the “deterministic” contribution from $t\approx 0$), we might conjecture that for any fixed $A>0$ and small $\varepsilon>0$, we have, very roughly speaking,
$\frac{1}{T}\int_{\begin{subarray}{c}|t|\leqslant T\\ t\not\approx 0\end{subarray}}\Bigl|\sum_{\begin{subarray}{c}n\leqslant x\\ P(n)\leqslant x^{\varepsilon}\end{subarray}}n^{-it}\Bigr|\,dt\ll\mathbb{E}\Bigl|\sum_{\begin{subarray}{c}n\leqslant x\\ P(n)\leqslant x^{\varepsilon}\end{subarray}}f(n)\Bigr|,\quad\forall x\leqslant T^{A}.$ (1.2)
Importantly, we know from Theorem 1.1 that the right-hand side is $o\bigl(\sqrt{\Psi(x,x^{\varepsilon})}\bigr)$. Similar to as described in [harpertypical], though isolating the main term in this case, Perron’s formula suggests that
$\frac{\Psi(x+h,x^{\varepsilon})-\Psi(x,x^{\varepsilon})}{h}\approx\frac{1}{2\pi x}\int_{-x/h}^{x/h}\sum_{\begin{subarray}{c}n\leqslant x+h\\ P(n)\leqslant x^{\varepsilon}\end{subarray}}n^{-it}\biggl(\frac{x}{h}\biggr)^{it}\,dt\approx\frac{\Psi(x,x^{\varepsilon})}{x}+E(x,x^{\varepsilon},h),$
where
$E(x,x^{\varepsilon},h)\ll\frac{1}{x}\int_{\begin{subarray}{c}|t|\leqslant T\\ t\not\approx 0\end{subarray}}\Bigl|\sum_{\begin{subarray}{c}n\leqslant x+h\\ P(n)\leqslant x^{\varepsilon}\end{subarray}}n^{-it}\Bigr|\,dt.$
In particular, since $\Psi(x,x^{\varepsilon})\asymp_{\varepsilon}x$, combining Theorem 1.1 and (1.2) (with $A=2$) gives the bound $E(x,x^{\varepsilon},\sqrt{x})=o\big(\sqrt{\Psi(x,x^{\varepsilon})/x}\big)=o(1)$, which (very roughly) gives an asymptotic for the count in intervals of length $h=\sqrt{x}$. In fact, our quantitative result for $y$ in this range, Theorem LABEL:t:small_u_upper_bound, would suggest further that this strategy could show the existence of $x^{\varepsilon}$-smooth numbers in intervals of length $h\gg\sqrt{x}/(\log\log x)^{1/4}$, similarly to how the analogous conjecture in [harpertypical] gives Möbius cancellation in sums of length $h\gg\sqrt{x}/(\log\log x)^{1/4}$.
Based on Theorem 1.1, it seems reasonable to conjecture that “better than squareroot cancellation” should hold uniformly for all $y\leqslant x$.
Conjecture 1.2.
Theorem 1.1 holds uniformly for any $2\leqslant y\leqslant x$.
It would be nice if one could solve Conjecture 1.2 completely, though our main theorem covers what we believe to be the most interesting range. Analogously to (1.1), we always work by connecting the partial sums to the Euler product. To illustrate how the analysis changes for smaller $y$, in the most extreme case, when $y=2$, we have
$\mathbb{E}\Bigl|\sum_{\begin{subarray}{c}n\leqslant x\\ P(n)\leqslant 2\end{subarray}}f(n)\Bigr|=\mathbb{E}\Bigl|\sum_{n\leqslant\lfloor\frac{\log x}{\log 2}\rfloor}f(2)^{n}\Bigr|=\int_{0}^{1}\Bigl|\sum_{n\leqslant\lfloor\frac{\log x}{\log x}\rfloor}\mathrm{e}(n\theta)\Bigr|\,d\theta\ll\log\log x,$
which is very small relative to $\Psi(x,2)^{1/2}\asymp(\log x)^{1/2}$. This provides evidence towards our conjecture, and also points out that for small $y$, the problem is related to finding cancellation in exponential sums over lattices. The exponential sum estimate becomes challenging once $y\gg(\log x)^{1+o(1)}$, and new ideas are needed to cover this range.
In this paper, we observe two distinct sources for the “better than squareroot cancellation” phenomena in Theorem 1.1; the size of $y$ being key to this analysis. First of all, for $y$ being close to $x$, one should expect that the critical GMC phenomenon uncovered by Harper [HarperLow] should give better than squareroot cancellation. We find this to be the case, and moreover, we are able to use critical GMC phenomenon to prove better than squareroot cancellation even when $y$ is as small as $e^{(\log\log x)^{2}}$, say. We outline how this is done in Section LABEL:s:_large_y_explained, and a precise statement is given in Theorem LABEL:t:small_u_upper_bound. The critical GMC saving in our case specifically comes from exploiting the randomness of $f(p)$ on the small primes (to be precise, primes smaller than roughly $e^{\frac{1}{1-\alpha}}$ for $\alpha=\alpha(x,y)$ the saddle point, see Section LABEL:s:smooths_intro for details), and, as one may expect, the saving is always quite small (like a power of $\log\log x$, or smaller).
Given the previous works [HarperLow, Xu, Caichshort], one may not expect to observe cancellation that is larger than some power of $\log\log x$ for any reasonably large $y$. However, we find that when $y$ is smaller than roughly $x^{\frac{1}{\log\log x}}$, we can obtain an unexpectedly large saving (actually coming from the primes closer to $y$). In short, this saving comes from the fact that the dominant contribution to certain Euler product expectations comes from highly unlikely events (we explain this in detail in Section LABEL:s:outline_of_moderate_y). The flavour of this might remind some readers of the supercritical GMC case, although this is not exactly what we observe, since our Euler products are shifted off the half-line and are no longer approximated by the exponential of a log-correlated Gaussian field. This novel phenomenon allows us to prove the following:
Theorem 1.3.
Let $u=\frac{\log x}{\log y}$ and fix $\varepsilon>0$. For $e^{(\log\log x)^{1+\varepsilon}}\leqslant y\leqslant x^{\frac{1}{(\log\log x)^{1+\varepsilon}}}$, we have
$\mathbb{E}\Bigl|\sum_{\begin{subarray}{c}n\leqslant x\\ P(n)\leqslant y\end{subarray}}f(n)\Bigr|\ll\sqrt{\Psi(x,y)}$} (1.3)